OpenStudy (ellecullen):

At a hot dog stand, there is a fixed price for hot dogs and a different price for sift drinks. 1 person bought 3 hot dogs and 2soft drinks for $9.25.Aperson ahead bought 2 hot dogs and 1 soft drink. How much does 1 hot dog cost? I need help understanding why the right answer is $2.25

OpenStudy (alexrobin13):

How much did the second purchase cost?

OpenStudy (ellecullen):

$5.75

OpenStudy (alexrobin13):

Okay so let x be hot dogs and r be soft drinks. 3x+2y=9.25 2x+1y=5.75

OpenStudy (alexrobin13):

To find how much a hot dog costs, you need to isolate y in each equation.

OpenStudy (alexrobin13):

So 3x+2y=9.25 2y=9.25-3x y=(9.25-3x)/2

OpenStudy (alexrobin13):

And 2x+y=5.75 y=5.75-x

OpenStudy (alexrobin13):

Since y=y, you can set the second half of both of these equations together

OpenStudy (alexrobin13):

So y=y (9.25-3x)/2=5.75-2x 9.25-3x=11.50-4x x=11.50-9.25 x=2.25 So there you go :)